# Self-Employment as a Signal: Career Concerns with Hidden Firm Performance

### v1

@@ Abstract
We analyze a dynamic labor market in which a worker with career concerns chooses each period between (i) self-employment that makes output publicly observable and (ii) employment at a firm that pays a flat wage but keeps individual performance hidden from outside observers. Output is binary and the worker is risk averse. The worker values future opportunities through a reputation for talent; firms may be benchmark (myopic) (ignoring the informational content of an application) or equilibrium (updating beliefs from the very act of applying). Three forces shape equilibrium: an insurance–information trade-off, selection by reputation, and inference from application decisions. We show that (i) an absorbing employment region exists in which low-reputation workers strictly prefer the firm’s insurance and optimally cease producing public information; (ii) sufficiently strong reputation triggers self-employment in order to generate public signals and preserve future outside options; and (iii) with equilibrium firms, application choices act as signals that shift hiring thresholds and wages even when in-firm performance remains opaque. Comparative statics deliver sharp, testable predictions for the prevalence of self-employment, the cyclicality of switching, and wage dynamics across markets with different degrees of performance transparency. The framework links classic career-concerns models to contemporary environments in which some tasks generate portable, public histories while firm tasks remain unobserved by the outside market (e.g., open-source contributions, freelancing platforms, or sales roles with standardized public metrics). Our results rationalize recent empirical findings on the value of public performance records and illuminate when opacity inside firms dampens or amplifies reputational incentives.
@@
career concerns , reputation , self-employment , dynamic signaling , adverse selection
@@ 1 Introduction
Markets often see some but not all of what workers do. [^1] Think open-source commits and package releases for developers, contest entries or preprints for researchers, portfolio pieces for designers, client reviews for freelancers, and seller ratings on platforms. A developer’s open-source commits, a freelancer’s platform ratings, or a salesperson’s client-visible metrics are public; most day-to-day contributions inside firms are not. We study a dynamic environment in which a risk-averse worker cares about both current consumption and a reputation for talent that shapes future wages. Each period the worker chooses between self-employment (publicly observed binary output) and hidden firm employment (a flat wage, with individual performance not disclosed to the outside market).222“Hidden” means the market does not see an informative individual-level signal. Inside the firm, managers may observe rich performance signals; we discuss partial public observability in Section 7. The market maintains a public belief about the worker’s talent summarized by a Beta posterior; self-employment updates this public state via observed successes or failures, while hidden employment leaves it unchanged. [^3] Conjugacy keeps the public state finite-dimensional. None of the comparative statics requires Beta priors; any regular one-dimensional prior with monotone likelihood ratio yields the same single-crossing logic. Firms post flat wages. In a benchmark (myopic) case, wages are set equal to the public posterior mean, ignoring the information in application decisions. In the equilibrium case, firms infer from who is expected to apply and post the mean talent of the implied applicant pool.
@@
This setting captures a common choice in modern labor markets: whether to invest in a public track record that travels across jobs, or to accept an opaque wage inside a firm. The trade-off is first-order. Self-employment exposes the worker to risk but creates portable information that improves future outside options; hidden employment insures current consumption but slows market learning. The paper’s contribution is to characterize the dynamic interaction of this insurance–information trade-off with selection into public production and with firms’ inference from applications, delivering sharp comparative statics and testable implications for occupations with portable public histories.
@@
The worker’s problem has a simple structure. Because current public production raises expected utility more for higher-ability workers and because the posted wage is type-independent in a period, the difference between choosing public production and hidden employment is strictly increasing in talent and strictly decreasing in the posted wage. We formalize this single-crossing property in Lemma 1.444The result survives multi-period wages that are independent of contemporaneous individual output and extends to strictly concave $u$ beyond our normalizations. The only knife-edge failure is when current wages do condition on individual output publicly, which restores public learning inside the firm. It implies cutoff policies in the public state and underpins all subsequent results.
Our first main result is an absorbing hidden-employment region. In the benchmark (myopic) case, the posted wage equals the public posterior mean. If a worker below the cutoff takes a hidden job, the public state does not change next period, so the same comparison repeats—hidden employment is absorbing (Proposition 2). The same logic extends to equilibrium posting, where wages account for application-based inference (Proposition 1). Intuition: for sufficiently low public beliefs, the insurance value of a flat wage dominates the informational value of producing another public signal, and because hidden employment generates no public updates, the choice becomes self-confirming.
@@
Our second result establishes existence and uniqueness of the equilibrium cutoff when firms infer from applications. Equilibrium posting ties the wage to acceptance: the posted wage equals the truncated-posterior mean of the types who apply, and the cutoff solves the fixed point. Existence and uniqueness follow from the monotonicity in Lemma 1: the applicant-pool mean is strictly increasing in the cutoff, while indifference is strictly decreasing in the wage (Proposition 4). Intuition: adding marginally stronger applicants raises the equilibrium wage, which makes employment relatively more attractive for slightly weaker types, pinning down a unique fixed point.
Our third result compares equilibrium to the benchmark. Because equilibrium wages price the selected applicant pool, they are (weakly) below benchmark wages at the same public state. Consequently, the equilibrium acceptance cutoff is (weakly) lower (Proposition 5). Intuition: inference from applications corrects the myopic benchmark for adverse selection, compressing wages and tightening self-selection.
@@
Roadmap. Section 2 reviews related work. Section 3 presents the model. Section 4 characterizes equilibrium under benchmark (myopic) and equilibrium firms and establishes the absorbing-employment result. Section 6 studies comparative statics and testable implications. Section 7 discusses extensions and empirical connections. Section 8 concludes.
@@ 2 Related Literature
Our analysis speaks to facts from modern labor markets. Online platforms and open-source ecosystems provide portable, public records of performance; causal evidence shows that public signals substantially improve workers’ subsequent outcomes (Pallais, 2014; Stanton and Thomas, 2016; Horton, 2017). At the same time, much productive activity is not externally visible, and employers gradually learn about workers through noisy, partly private signals (Altonji and Pierret, 2001). By allowing workers to choose between public and hidden production technologies, our framework offers a unified explanation for when we should observe self-employment as a reputational strategy and when we should see long spells of opaque employment with limited outside learning.
@@
We connect classic career-concerns models (Holmström, 1999; Dewatripont et al., 1999a, b; Gibbons and Murphy, 1992) to environments with coexisting public and hidden performance. Relative to that literature, we (i) endogenize the observability choice by letting workers switch between self-employment and hidden firm employment; (ii) characterize an absorbing employment region that rationalizes persistent opacity and slow social learning; and (iii) show how application decisions themselves become signals for equilibrium firms. These elements yield predictions for the incidence and timing of switching, wage dynamics across transparency regimes, and the response to changes in risk exposure or output variance.
@@
Our starting point is the dynamic-career-concerns framework in which market beliefs about talent respond to observed performance (Holmström, 1999). The structure and publicity of information are central to incentives: better information can increase or decrease effort depending on how signals reach the market and how they are aggregated (Dewatripont et al., 1999a, b). We also draw on Gibbons and Murphy (1992), who show that explicit incentive contracts should optimally account for the implicit incentives coming from career concerns. In our setting, the implicit incentives arise predominantly from the worker’s choice of where to produce information (publicly vs. opaquely), rather than from an explicit pay-for-performance scheme.
@@
Our analysis is related to dynamic reputation models in which current actions affect future beliefs (e.g., Board and Meyer-ter-Vehn, 2013). We differ by making the observability of outcomes endogenous through the employment-mode decision. This generates selection on reputation: relatively strong types choose public production to separate, which amplifies the informativeness of public signals; weak types value insurance and hide. The result is an absorbing employment region in which workers optimally stop producing public information—a mechanism that slows social learning even in otherwise transparent markets.
@@
A second thread studies the social value of early-career information and the difficulty firms face in appropriating its benefits. Terviö (2009) shows how markets can under-invest in talent discovery when the returns to information about novice workers are largely external. Empirically, making performance public on online labor markets or via credible third parties causally improves workers’ subsequent outcomes (Pallais, 2014; Stanton and Thomas, 2016; Horton, 2017). Our theory provides conditions under which workers voluntarily seek such transparency by going solo, and conditions under which they rationally prefer opaque employment.
@@
A large empirical literature documents how employers learn about worker productivity over time and the extent to which that learning is public or private (Altonji and Pierret, 2001). We contribute a distinct mechanism: when firm performance is not disclosed externally, the act of applying for such employment becomes informative to equilibrium employers. This echoes broader insights that markets learn not only from realized outcomes but also from selection and participation decisions. In equilibrium, wages and hiring thresholds internalize the informational content of applications even if post-hire outputs remain opaque.
@@
Our applications to open-source contributions and freelancing platforms connect to research on reputation systems and public track records (Bar-Isaac and Tadelis, 2008). In such settings, outside audiences can readily observe some outputs, while many internal tasks remain hidden. The predictions of our model—about switching, absorbing hidden employment for weaker reputations, and wage dynamics across transparency regimes—organize a number of recent empirical facts and suggest new tests, such as how shocks to the visibility of output (e.g., the introduction of new public metrics) shift the incidence of self-employment and the slope of wage–reputation profiles across markets.
Across these strands, our contribution is to endogenize the generation of public information through an employment-mode choice and to analyze how this interacts with firm sophistication in inference. Doing so yields new comparative statics and testable predictions for the distribution of public track records, the clustering of opaque employment spells, and the mapping from reputation to wages when some production is deliberately hidden.
@@ 3 Model
Environment
Time is discrete, $t=0,1,2,\ldots$, and the common discount factor is $\delta\in(0,1)$. A single worker has a fixed talent $\theta\in[0,1]$ that is known to the worker and unknown to the outside market. [^5] Allowing for worker learning about own talent is discussed in an extension; here talent is fixed and privately known. In each period the worker chooses between (i) self-employment ($S$), which yields a binary outcome $y_{t}\in\{0,1\}$ with $\Pr(y_{t}=1\mid\theta)=\theta$ that is publicly observed by the market, and (ii) employment at a firm ($E$), which pays a flat wage $w_{t}$ but keeps the worker’s individual outcome hidden from the outside market. In our baseline, the firm does not disclose individual outcomes externally. [^6] In Section 7 we allow the firm to observe a private signal and discuss renegotiation.
Preferences and payoffs
Consumption in period $t$ equals $c_{t}=y_{t}$ under $S$ and $c_{t}=w_{t}$ under $E$ (we normalize the monetary payoff of a success to $1$). The worker has Bernoulli utility $u:\mathbb{R}_{+}\to\mathbb{R}$ that is increasing and concave, with normalization $u(0)=0$ and u(1)=1u(1)=1777This pins units so the expected period utility from self-employment equals the success probability $\theta$. Any positive affine transformation of $u$ would merely re-scale wages; all choices are invariant.. Hence the current utility from $S$ is $\theta$ and from $E$ is $u(w_{t})$. Risk aversion (concavity of $u$) matters through comparison with the certain wage under $E$ and through the value of reputation for future wages.
Beliefs and the public state
@@
A period of hidden employment leaves $(\alpha,\beta)$ unchanged. We write $\mu_{t}\equiv\mu(\alpha_{t},\beta_{t})$.
Firms and wage posting
Each period a competitive fringe of short-lived firms posts a flat wage for applicants. Firms cannot condition the wage on contemporaneous individual performance and do not disclose worker-specific outcomes to the market. [^9] Think unverifiable or multi-task output, team production, or high disclosure costs. Allowing the firm to privately observe a signal and renegotiate does not restore public learning; the absorbing logic only vanishes when the market receives a sufficiently informative public signal from firm employment (Section 7). A firm that hires a worker of talent $\theta$ produces expected revenue $\theta$ (equal to the success probability) and pays wage $w_{t}$. Competition and free entry imply zero expected profits, so the posted wage equals the expected talent of the applicant pool relevant to that firm in that period.
@@
  • Equilibrium firms. Firms correctly infer that application/acceptance is endogenous and update from the application decision. Let $\Theta^{E}(\alpha,\beta)\subseteq[0,1]$ denote the set of types who apply at state $(\alpha,\beta)$ under the worker’s strategy. Then posted wages satisfy
    $w^{S}(\alpha,\beta)\;=\;\mathbb{E}\big{[}\theta\mid\theta\in\Theta^{E}(\alpha,\beta);\;\alpha,\beta\big{]},$ (2)
[…]
In equilibrium under mild regularity (proved later), the worker’s strategy is monotone in $\theta$ and $\Theta^{E}(\alpha,\beta)$ is an interval $0,\hat{\theta}(\alpha,\beta)]$; then $w^{S}$ is the truncated-posterior mean.
Timing within a period
@@
The worker’s problem
Let $V(\theta;\alpha,\beta)$ denote the worker’s value function. For a given wage policy $w(\alpha,\beta)$ (naïve or equilibrium), the Bellman equation is
@@
Single crossing and cutoff strategies
Define $\Delta(\theta;\alpha,\beta)\equiv U_{S}(\theta;\alpha,\beta)-U_{E}(\theta;\alpha,\beta)$. Under (3)–(5), $\Delta(\theta;\alpha,\beta)$ is strictly increasing in $\theta$: higher $\theta$ raises current expected utility under $S$ and increases the probability of transitioning to the higher-mean state $(\alpha+1,\beta)$. [^11] Formally, $U_{S}$ is strictly increasing in $\theta$, while $U_{E}$ depends on $\theta$ only through the continuation value $V(\theta;\alpha,\beta)$, which is increasing but with slope bounded by $\delta$; the current term in $U_{S}$ delivers the strict single crossing. A complete proof is in Appendix A. Hence there exists a unique cutoff $\hat{\theta}(\alpha,\beta)\in[0,1]$ such that types $\theta\leq\hat{\theta}(\alpha,\beta)$ prefer employment and types $\theta>\hat{\theta}(\alpha,\beta)$ prefer self-employment:
@@
In the equilibrium case, conditions (1) and (2) jointly pin down the fixed point $(\hat{\theta}(\alpha,\beta),w^{S}(\alpha,\beta))$.
Absorbing employment region
@@
Intuitively, for low-reputation public states $(\alpha,\beta)$, the combination of (i) a low posted wage (especially when firms are equilibrium and price the selected applicant pool) and (ii) risk aversion makes the certain wage comparatively attractive for lower types. Because hidden employment produces no new public information, the belief state does not improve, and the same comparison repeats indefinitely. [^12] Formally, the public state $(\alpha,\beta)$ is a controlled Markov state that only moves under self-employment; choosing hidden employment is a self-confirming policy at any state where it is optimal.
Benchmark versus equilibrium pricing
In the naïve benchmark, wages equal the public mean $\mu(\alpha,\beta)$ regardless of application. In the equilibrium case, the wage equals the mean of the truncated posterior on $0,\hat{\theta}(\alpha,\beta)]$. The two benchmarks therefore generate different cutoffs: the equilibrium wage $w^{S}$ incorporates the adverse selection induced by (6), which generally lowers $w$ relative to $w^{N}$ and feeds back into the worker’s choice. Section 4 characterizes the fixed point and shows how pricing sophistication shifts $\hat{\theta}(\alpha,\beta)$ and the size of the absorbing region.

### v3

@@ Abstract
We study a dynamic labor market in which a risk-averse worker with career concerns chooses each period between self-employment, which generates publicly observed binary output, and employment at a firm, which pays a flat wage but keeps individual performance hidden from the outside market. The worker values future opportunities through a reputation for talent, understood as the public belief used by firms to set wages. Firms may be myopic, pricing only off public beliefs, or sophisticated, inferring from who is expected to apply. Three forces organize the results: an insurance–information trade-off, selection by talent, and inference from application decisions. A self-confirming absorbing hidden-employment region arises naturally when the value of insurance dominates the value of additional public signals, so some workers optimally stop generating public information. At any public belief there is a unique talent cutoff governing selection into self-employment, and sophisticated firms, by internalizing application-based inference, compress wages and induce different patterns of public information production than myopic firms. The framework yields testable predictions for the prevalence of self-employment, switching into opaque jobs, and wage dynamics across markets with different degrees of performance transparency.
@@
Career concerns , reputation , signaling , self-employment , transparency , labor-market learning
@@ 1 Introduction
Markets often see some but not all of what workers do. A developer’s open-source commits, a freelancer’s platform ratings, or a salesperson’s client-visible metrics are public; most day-to-day contributions inside firms are not. We study a dynamic environment in which a risk-averse worker cares about both current consumption and a reputation for talent that shapes future wages. Each period the worker chooses between self-employment (publicly observed binary output) and hidden222“Hidden” means the market does not see an informative individual‐level signal. Inside the firm, managers may observe rich performance signals; we discuss partial public observability in Section 6. firm employment (a flat wage, with individual performance not disclosed to the outside market). The market maintains a public belief about the worker’s talent summarized by a Beta posterior; self-employment updates this public state via observed successes or failures, while hidden employment leaves it unchanged. [^3] Conjugacy keeps the public state finite‐dimensional. None of the comparative statics requires Beta priors; any regular one‐dimensional prior with monotone likelihood ratio yields the same single‐crossing logic. Firms post flat wages. In a benchmark case, wages are set equal to the public posterior mean, ignoring the information in application decisions. When firms are sophisticated, they infer from who is expected to apply and post the mean talent of the implied applicant pool.
We embed this choice in a stationary, infinite-horizon environment with overlapping generations of workers. Time is discrete and there is a constant probability that any given worker exits the market in each period; exiting workers are replaced by new entrants drawn from the same prior. Firms and outside employers do not observe calendar age or total tenure, only the worker’s public record of self-employment outcomes. In equilibrium, wages and cutoff strategies are therefore time-invariant functions of the public state, and our comparative statics can be interpreted as describing steady-state differences across labor markets. [^4] This infinite-horizon, overlapping-generations formulation follows the spirit of Pei, (2025)’s analysis of stationary reputational environments.
@@
This setting captures a common choice in modern labor markets: whether to invest in a public track record that travels across jobs, or to accept an opaque wage inside a firm. Equivalently, one can read the model as a worker deciding how much effort to allocate to public projects—open-source contributions, contests, portfolio pieces, client-visible work—versus hidden projects inside the firm whose outcomes are not externally observable. The trade-off is first-order. Self-employment exposes the worker to risk but creates portable information that improves future outside options; hidden employment insures current consumption but slows market learning. The paper’s contribution is to show how selection into self-employment versus hidden jobs jointly shapes both wages and the amount of public information produced about talent, under naïve and sophisticated wage-setting regimes, and to characterize the resulting comparative statics and testable implications for occupations with portable public histories.
@@
The worker’s problem has a simple structure. Because current public production raises expected utility more for higher-ability workers and because the posted wage is type-independent in a period, the difference between choosing public production and hidden employment is strictly increasing in talent and strictly decreasing in the posted wage. We formalize this single-crossing property in Lemma 1.555The result survives multi‐period wages that are independent of contemporaneous individual output and extends to strictly concave $u$ beyond our normalizations. The only knife‐edge failure is when current wages do condition on individual output publicly, which restores public learning inside the firm. It implies cutoff policies in the public state and underpins all subsequent results.
First, we characterize a self-confirming absorbing hidden-employment region in the state space. For sufficiently low public beliefs about the worker’s talent, the flat wage offered by firms dominates the option value of generating further public signals, so the worker optimally chooses hidden employment; because that choice produces no public updates, the worker then remains in this region going forward. The size and shape of this region respond in intuitive ways to primitives such as risk aversion, patience, and the informativeness of public performance: it expands when workers are more risk averse or opaque wages are generous, and shrinks when workers are more patient, better insured, or operate in more transparent labor markets.
@@
A second key implication is the existence and uniqueness of a well-behaved cutoff: at each public state there is a single talent threshold separating workers who self-employ (to generate or maintain public signals) from those who choose the firm’s insurance, and standard monotonicity arguments ensure that this fixed point is uniquely pinned down. Existence and uniqueness follow from the monotonicity in Lemma 1: the applicant-pool mean is strictly increasing in the cutoff, while indifference is strictly decreasing in the wage (Proposition 4). Intuition: adding marginally stronger applicants raises the equilibrium wage, which makes employment relatively more attractive for slightly weaker types, pinning down a unique fixed point.
We also compare sophisticated pricing to the myopic benchmark: because sophisticated firms internalize adverse selection from application decisions, they offer lower wages than naïve firms, compress the payoff to additional public signals, and tighten self-selection into self-employment versus hidden jobs. Methodologically, our stationary infinite-horizon formulation is in the spirit of Pei, (2025), who studies reputational dynamics with endogenous records in a doubly infinite horizon, and conceptually our hidden-effort extension connects the worker’s observability choice to the work–shirk dynamics in dynamic reputation models such as Board and Meyer-ter Vehn, (2013).
@@
Roadmap. Section 2 reviews related work. Section 3 presents the model. Section 4 characterizes equilibrium under myopic and sophisticated wage setting and establishes the absorbing-employment result. Section 5 develops comparative statics and empirical implications. Section 6 discusses extensions, policy implications, and alternative interpretations. Section 7 concludes.
@@ 2 Related Literature
Our analysis speaks to facts from modern labor markets. Online platforms and open-source ecosystems provide portable, public records of performance; causal evidence shows that public signals substantially improve workers’ subsequent outcomes (Pallais,, 2014; Stanton and Thomas,, 2016; Horton,, 2017). At the same time, much productive activity is not externally visible, and employers gradually learn about workers through noisy, partly private signals (Altonji and Pierret,, 2001). By allowing workers to choose between public and hidden production technologies, our framework offers a unified explanation for when we should observe self-employment as a reputational strategy and when we should see long spells of opaque employment with limited outside learning.
@@
We connect classic career-concerns models (Holmström,, 1999; Dewatripont et al., 1999a, ; Dewatripont et al., 1999b, ; Gibbons and Murphy,, 1992) to environments with coexisting public and hidden performance. Relative to that literature, we (i) endogenize the observability choice by letting workers switch between self-employment and hidden firm employment; (ii) characterize an absorbing employment region that rationalizes persistent opacity and slow social learning; and (iii) show how application decisions themselves become signals for sophisticated firms. These elements yield predictions for the incidence and timing of switching, wage dynamics across transparency regimes, and the response to changes in risk exposure or output variance.
@@
Our starting point is the dynamic-career-concerns framework in which market beliefs about talent respond to observed performance (Holmström,, 1999). The structure and publicity of information are central to incentives: better information can increase or decrease effort depending on how signals reach the market and how they are aggregated (Dewatripont et al., 1999a, ; Dewatripont et al., 1999b, ). We also draw on Gibbons and Murphy, (1992), who show that explicit incentive contracts should optimally account for the implicit incentives coming from career concerns. In our setting, the implicit incentives arise predominantly from the worker’s choice of where to produce information (publicly vs. opaquely), rather than from an explicit pay-for-performance scheme.
@@
Our analysis is related to dynamic reputation models in which current actions affect future beliefs (e.g., Board and Meyer-ter Vehn,, 2013). We differ by making the observability of outcomes endogenous through the employment-mode decision. This generates selection on reputation: relatively strong types choose public production to separate, which amplifies the informativeness of public signals; weak types value insurance and hide. The result is an absorbing employment region in which workers optimally stop producing public information—a mechanism that slows social learning even in otherwise transparent markets.
More broadly, our model is close in spirit to recent work on dynamic career concerns with rich information environments. (Pei,, 2025) analyzes community enforcement with endogenous records in anonymous matching: players can add or erase signals from their histories, and the resulting monotonicity constraints on continuation values sharply restrict cooperation when lifetimes are very long. Our worker’s choice between generating new public signals and freezing the record through hidden employment is a labor-market analogue of this record-design margin. Other recent dynamic signaling models with career concerns emphasize experimentation and complex information flows (e.g. Thomas,, 2019; Kaniel and Orlov,, 2025). On the empirical side, Abou El-Komboz and Goldbeck, (2025) show that developers’ career concerns causally increase open-source contributions, while Belenzon and Schankerman, (2015) document sorting and signaling motives in open innovation. These findings motivate our focus on self-employment and public projects as a reputational technology and on how market transparency shapes the supply of public information about talent.
A large literature studies dynamic careers in internal labor markets. In Gibbons and Waldman, (1999), firms learn about worker ability from internal output and adjust job assignments and wages over time; Farber and Gibbons, (1996) develop a dynamic model of learning and wage dynamics in a competitive labor market; and Gibbons and Waldman, (2006); Waldman, (2012) survey theory and evidence on careers in organizations. In these models, performance is effectively observed by the relevant decision maker (the firm), so learning is primarily about who to promote or retain. By contrast, we focus on environments where some production is deliberately hidden from the outside market, so that the worker’s choice between public and hidden work jointly determines both the information set of future employers and the mapping from reputation into wages. Our absorbing hidden-employment region can be read as a simple mechanism through which careers become opaque and social learning slows down, even when performance is transparent inside firms.
Related empirical work studies learning and selection in online labor markets and digital platforms. Leung, (2018) shows how employers update beliefs about worker quality from their own past hiring experiences in a large online freelance market, while Kokkodis and Ipeirotis, (2014) document how platform ratings and skill signals shape demand and match quality. These papers emphasize that observed performance and feedback in digital environments are central inputs into hiring and pricing decisions. Our framework complements this evidence by modeling a worker’s choice to invest in public versus hidden performance as a reputational technology, and by characterizing how different wage-setting regimes (naïve versus sophisticated) affect both selection into public work and the amount of public information generated about talent.
@@
A second thread studies the social value of early-career information and the difficulty firms face in appropriating its benefits. Terviö, (2009) shows how markets can under-invest in talent discovery when the returns to information about novice workers are largely external. Empirically, making performance public on online labor markets or via credible third parties causally improves workers’ subsequent outcomes (Pallais,, 2014; Stanton and Thomas,, 2016; Horton,, 2017). Our theory provides conditions under which workers voluntarily seek such transparency by going solo, and conditions under which they rationally prefer opaque employment.
@@
A large empirical literature documents how employers learn about worker productivity over time and the extent to which that learning is public or private (Altonji and Pierret,, 2001). We contribute a distinct mechanism: when firm performance is not disclosed externally, the act of applying for such employment becomes informative to sophisticated employers. This echoes broader insights that markets learn not only from realized outcomes but also from selection and participation decisions. In equilibrium, wages and hiring thresholds internalize the informational selection into opaque jobs versus continued public production, and in our sophisticated pricing case wages are set using the expected applicant pool even when post-hire outputs remain opaque.
@@
Our applications to open-source contributions and freelancing platforms connect to research on reputation systems and public track records (Bar-Isaac and Tadelis,, 2008). In such settings, outside audiences can readily observe some outputs, while many internal tasks remain hidden. The predictions of our model—about switching, absorbing hidden employment for weaker reputations, and wage dynamics across transparency regimes—organize a number of recent empirical facts and suggest new tests, such as how shocks to the visibility of output (e.g., the introduction of new public metrics) shift the incidence of self-employment and the slope of wage–reputation profiles across markets.
Across these strands, our contribution is to endogenize the worker’s choice between public and hidden production, to characterize the absorbing hidden-employment region in which workers rationally remain opaque, and to compare naïve and sophisticated wage-setting regimes that differ in how they map public reputation into wages when some production is deliberately hidden.
@@ 3 Model
Time is discrete, $t=0,1,2,\dots$, and the common discount factor is $\delta\in(0,1)$. A single worker has a fixed talent $\theta\in[0,1]$ that is known to the worker and unknown to the outside market. [^6] Formally, one can embed this “single worker” problem in a stationary overlapping-generations environment with a constant exogenous exit probability $\lambda\in(0,1)$ each period. The worker’s Bellman equation then has the continuation term multiplied by $\delta(1-\lambda)$ rather than $\delta$, so all of our results go through with an effective discount factor that incorporates both time preference and expected career length. We adopt this reduced-form formulation in the text for notational simplicity. Firms and outside employers do not observe the worker’s calendar age or total tenure; they only see the history of public outcomes generated under self-employment, which we later summarize by the Beta posterior parameters $(\alpha,\beta)$.
In each period the worker chooses between (i) self-employment ($S$), which yields a binary outcome $y_{t}\in\{0,1\}$ with $\Pr(y_{t}=1\mid\theta)=\theta$ that is publicly observed by the market, and (ii) employment at a firm ($E$), which pays a flat wage $w_{t}$ but keeps the worker’s individual outcome hidden from the outside market. In our baseline, the firm does not disclose individual outcomes externally. [^7] In Section 6 we allow the firm to observe a private signal and discuss renegotiation.
Consumption in period $t$ equals $c_{t}=y_{t}$ under $S$ and $c_{t}=w_{t}$ under $E$ (we normalize the monetary payoff of a success to $1$). The worker has Bernoulli utility $u:\mathbb{R}_{+}\to\mathbb{R}$ that is increasing and concave, with normalization $u(0)=0$ and u(1)=1u(1)=1888This pins units so the expected period utility from self‐employment equals the success probability $\theta$. Any positive affine transformation of $u$ would merely re‐scale wages; all choices are invariant.. Hence the current utility from $S$ is $\theta$ and from $E$ is $u(w_{t})$. Risk aversion (concavity of $u$) matters through comparison with the certain wage under $E$ and through the value of reputation for future wages.
@@
A period of hidden employment leaves $(\alpha,\beta)$ unchanged.
We write $\mu_{t}\equiv\mu(\alpha_{t},\beta_{t})$.
Each period a competitive fringe of short-lived firms posts a flat wage for applicants. Firms cannot condition the wage on contemporaneous individual performance and do not disclose worker-specific outcomes to the market. [^10] Think unverifiable or multi‐task output, team production, or high disclosure costs. Allowing the firm to privately observe a signal and renegotiate does not restore public learning; the absorbing logic only vanishes when the market receives a sufficiently informative public signal from firm employment (Section 6). A firm that hires a worker of talent $\theta$ produces expected revenue $\theta$ (equal to the success probability) and pays wage $w_{t}$. Competition and free entry imply zero expected profits, so the posted wage equals the expected talent of the applicant pool relevant to that firm in that period.
@@
  • Sophisticated firms. Firms correctly infer that application/acceptance is endogenous and update from the application decision. Let $\Theta^{E}(\alpha,\beta)\subseteq[0,1]$ denote the set of types who apply at state $(\alpha,\beta)$ under the worker’s strategy. Then posted wages satisfy
    $w^{S}(\alpha,\beta)=\mathbb{E}[\theta\mid\theta\in\Theta^{E}(\alpha,\beta);\alpha,\beta],$ (2)
    with the expectation taken under the $\mathrm{Beta}(\alpha,\beta)$ posterior. [^11] With a continuum of firms and applicants, the posted wage is the competitive price for the marginal applicant; with anonymous jobs and free entry, this equals the pool mean talent.
Throughout, we use the term reputation to refer to the public belief about the worker’s talent that is encoded in the posterior distribution and, in particular, in its mean $\mu(\alpha,\beta)$. This public belief is what firms use to set wages in both the naïve and sophisticated cases. In this sense, a worker’s “career concerns” are concerns about how current choices affect future realizations of $\mu(\alpha,\beta)$ and thus future wage offers.
In equilibrium under mild regularity, the worker’s strategy is monotone in $\theta$ and $\Theta^{E}(\alpha,\beta)$ is an interval $0,\hat{\theta}(\alpha,\beta)]$; then $w^{S}$ is the truncated-posterior mean.
@@
Let $V(\theta;\alpha,\beta)$ denote the worker’s value function. For a given wage policy $w(\alpha,\beta)$ (naïve or sophisticated), the Bellman equation is
@@
Define
$\Delta(\theta;\alpha,\beta)\equiv U_{S}(\theta;\alpha,\beta)-U_{E}(\theta;\alpha,\beta).$
Under (3)–(5), $\Delta(\theta;\alpha,\beta)$ is strictly increasing in $\theta$: higher $\theta$ raises current expected utility under $S$ and increases the probability of transitioning to the higher-mean state $(\alpha+1,\beta)$. [^12] Formally, $U_{S}$ is strictly increasing in $\theta$, while $U_{E}$ depends on $\theta$ only through the continuation value $V(\theta;\alpha,\beta)$, which is increasing but with slope bounded by $\delta$; the current term in $U_{S}$ delivers the strict single crossing. Hence there exists a unique cutoff $\hat{\theta}(\alpha,\beta)\in[0,1]$ such that types $\theta\leq\hat{\theta}(\alpha,\beta)$ prefer employment and types $\theta>\hat{\theta}(\alpha,\beta)$ prefer self-employment:
@@
In the sophisticated pricing regime, the worker’s cutoff condition (6) and the wage equation (2) jointly pin down the fixed point $(\hat{\theta}(\alpha,\beta),w^{S}(\alpha,\beta))$.
Given the infinite-horizon, overlapping-generations interpretation in Section 3, we focus on stationary Markov equilibria in which the worker’s strategy and wage functions depend only on the current public state $(\alpha,\beta)$ and, in the sophisticated case, on the implied applicant pool. The equilibrium characterized above is stationary in this sense: $(\sigma,w^{N},w^{S})$ do not depend on calendar time. Together with the constant inflow of new workers with prior $(\alpha_{0},\beta_{0})$ and the law of motion for beliefs, this induces a steady-state cross-sectional distribution of public states in the population. Our comparative statics can therefore be interpreted as describing how the steady-state allocation of workers between public and hidden employment responds to changes in primitives such as risk aversion, patience, and performance transparency.
@@
Intuitively, for low-reputation public states $(\alpha,\beta)$ the combination of (i) a low posted wage (especially when firms are sophisticated and price the selected applicant pool) and (ii) risk aversion makes the certain wage comparatively attractive for lower types. Because hidden employment produces no new public information, the public belief $(\alpha,\beta)$ does not improve, and the same comparison repeats in every subsequent period. Proposition 1 therefore implies a simple geometry for the absorbing hidden-employment region: at any public state $(\alpha,\beta)$ the set of types who eventually remain in employment forever is downward closed in talent and can be written as
$\mathcal{A}(\alpha,\beta)\;=\;\left\{\,(\alpha,\beta,\theta):\theta\leq\hat{\theta}(\alpha,\beta)\,\right\},$
with all $\theta\leq\hat{\theta}(\alpha,\beta)$ optimally choosing $E$ and generating no further public signals, while all $\theta>\hat{\theta}(\alpha,\beta)$ continue to self-employ to keep the option value of information alive. Along the path of a given true talent $\theta$, entry into this absorbing region can occur only after a sufficiently adverse sequence of public outcomes drives the public belief $\mu(\alpha,\beta)$ low enough that the insurance value of the flat wage outweighs the option value of another signal. Later, Proposition 5 shows that under sophisticated firms the cutoff satisfies $\hat{\theta}^{S}(\alpha,\beta)\leq\hat{\theta}^{N}(\alpha,\beta)$, so the absorbing region in $(\alpha,\beta,\theta)$-space is weakly smaller when firms fully internalize application-based inference. [^13] Formally, the public state $(\alpha,\beta)$ is a controlled Markov state. Choosing $E$ induces the deterministic transition $(\alpha,\beta)\to(\alpha,\beta)$, so any policy that prescribes $E$ for all $\theta\leq\hat{\theta}(\alpha,\beta)$ is self-confirming at a state where it is optimal.
In the naïve benchmark, wages equal the public mean $\mu(\alpha,\beta)$ regardless of application. In the sophisticated pricing regime, the wage equals the mean of the truncated posterior on $0,\hat{\theta}(\alpha,\beta)]$. The two benchmarks therefore generate different cutoffs: the sophisticated wage $w^{S}$ incorporates the adverse selection induced by (6), which generally lowers $w$ relative to $w^{N}$ and feeds back into the worker’s choice. Section 4 characterizes the fixed point and shows how pricing sophistication shifts $\hat{\theta}(\alpha,\beta)$ and the size of the absorbing region.